Observers: Walth, Penner, Clement

mostly clear, seeing ~1.5" (guide star)

19:10 telluric star, GCS=1.35"

20:07 done with telluric for Greg's masks, moving to Kyle's flux std

20:30 1.02" dimm 1.2" guider, thin cirrus

20:38 moving to A611 for science (Greg)

21:01 starting exposure in zJ, avg seeing 0.77" (guider), frame #58 to #63 (1 hour)

22:12 moving to A851 (Greg)

22:32 starting exposure in H, avg seeing guider 1.4"

00:43 moving to GOODS-N, will check seeing once at that position

01:36 seeing 1.5"-2"

01:48 seeing is still >1.5" + thin cirrus, moving to Ian's backup target (LUCI imaging).

01:55 starting imaging of J1414 in J band, first frame is #102,

02:55 #117-124 probably affected by clouds

03:05 thicker clouds coming in. Reflections seen on frames #126/127/128. Not guiding for #129, #130, #131.

03:29 resumed imaging on J1414 in J band (#132). Guider is ~1.2", varying.

04:13 last frame for J1414 is #146, frames after #135 probably affected by clouds

04:15 moving to MACS1149 (greg)

04:35 start integrating in K band (first #151), guider is 1.1

05:44 observations completed for MACS1149 (last #174), guider was 1.1-1.3", clouds

06:14 closing

-- Main.Clement - 05 Mar 2013
Topic revision: r7 - 05 Mar 2013, XiaohuiFan
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback